# Rice-like complexity lower bounds for Boolean and uniform automata networks

### v1

@@ 3.2 Metareduction
The neutral is intending to “change nothing”, and is obtained by pumping (see next Subsection). However, tools from finite model theory provide four graphs, meeting the requirements stated below.
Proposition 4 ([10, Proposition 19]).
Let $\psi$ be an MSO formula and $\Gamma=\{G_{0},G_{1},G_{2},G_{3}\}$ four $k$-graphs for some $k\in\mathbb{N}$, such that:
  1. 1.
    $|G_{0}|=|G_{1}|$,
  2. 2.
    $P_{1}(G_{0})\cap P_{2}(G_{0})=P_{1}(G_{1})\cap P_{2}(G_{1})\neq V(G_{1})$, and
  3. 3.
    for every $p$ in $P_{1}(G_{0})\cap P_{2}(G_{0})$, we have $G_{0}(p)=G_{1}(p)$.
@@
Suppose that, for every word $w$ over alphabet $\{0,1\}$, we have $\Delta^{\Gamma}(2\cdot w\cdot 3)\vDash\psi$ if and only if $w$ contains letter 00. Then $\psi$-NAN-dynamics is ${\mathsf{NP}}$-hard.
To prove the proposition, it requires to take into account vertices merged by the gluing operations, in order to construct a NAN $f$ of appropriate size, so that each configuration (vertex of $\mathcal{G}_{f}$) plays its expected role. We denote $\overline{S}$ the word of length $2^{s}$ on alphabet $\{0,1\}$ such that $\overline{S}_{i}$ is $1$ when the assignment $i$ satisfies $S$, and is 00 otherwise ($i$ being interpreted as a $s$ bits assignment). The reduction builds the circuit of a NAN whose dynamics is $\Delta^{\Gamma}(2\cdot\bar{S}\cdot 3)$. The only non-constant part of the construction consists in evaluating $S$.
3.3 Pumping models and counter-models
Let us sketch how the four graphs from Proposition 4 are obtained. In particular the neutral element, in the next proposition.
Proposition 5 ([10, Proposition 8]).
Let $\psi$ be an MSO formula and an integer $k$. If $\psi$ has infinitely many models of treewidth $k$, then there exist $\tilde{\Gamma}=\{\tilde{G_{1}},\tilde{G_{2}},\tilde{G_{3}}\}$ three $(k+1)$-graphs such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)\vDash\psi$ for all $\ell\in\mathbb{N}$, and $P_{1}(\tilde{G_{1}})\cap P_{2}(\tilde{G_{1}})\neq V(G_{1})$.
The proof relies on Remark 2. Since $\Sigma_{k}$ is finite, considering a sufficiently big model ensures that there exists a part that can be pumped, denoted $\tilde{G_{1}}$. It is identified on tree-decompositions, and glued in terms of graphs thanks to Remark 2.
In order to furhtermore obtain Conditions 1–3 of Proposition 4, the graph $\Omega_{m}$ will play the role of $G_{0}$. If $\Omega_{m}$ is a counter-model of $\psi$, then we consider $\neg\psi$ instead of $\psi$ (so that $\Omega_{m}$ is sufficient; and Proposition 5 is applied on $\neg\psi$). Observe that $\Omega_{m}$ can accomodate any disjoint union yet remain saturating. Take $G_{0}$ as the disjoint union of $\Omega_{m}$ and $\tilde{G_{1}}$. On the other side, take $G_{1}=\bigoplus_{i=1}^{\alpha}\tilde{G_{1}}$ with $\alpha$ such that $|G_{1}|\geq|G_{0}|$. Finaly, pad $G_{0}$ in order to have $|G_{0}|=|G_{1}|$. Then with $G_{2}=\tilde{G_{2}}$ and $G_{3}=\tilde{G_{3}}$ we indeed obtain all the resquirements for Proposition 4.
To conclude the hardness proof by reduction from SAT or UNSAT, it remains to construct the circuit of an automata network such that the dynamics has: one copy of $G_{2}$, then copies of $G_{0}$ or $G_{1}$ for the $2^{\ell}$ subsets of configurations on which the formula $S$ is evaluated, and finaly one copy of $G_{3}$. All these copies also need to be glued correctly by the circuit, which is easy to process in polynomial time.
@@ 3.4 Contributions
  • A careful consideration in gluing deterministic dynamics (graphs of out-degree $1$ are handled through Hanf’s local lemma in [9]) in Section 4.
  • The adjunction of arithmetical considerations in Section 5, to restrict the constructions to $q$-uniform networks in Section 6.
@@ 4 Complexity lower bounds for deterministic networks
In this section, we extend Theorem 1.a for FO from FO to MSO logics, or more accurately we transpose the proof technique of Theorem 1.c from non-deterministic to deterministic dynamics, which are succinctly represented graphs of out-degree $1$. Recall that the non-triviality of formulas is defined relative to graphs of out-degree $1$ only. We prove the following part of Theorem 1.
@@
We follow the proof structure presented in Section 3, using the deterministic saturating graph from Corollary 3. The main difficulty consists in transposing Proposition 5 in order to obtain graphs of out-degree $1$ all along the way (i.e., to ensure that gluings preserve this property), and meet Conditions 1–3 of Proposition 4. We combine these into the following proposition (out-degree $1$ graphs have treewidth at most $2$).
@@
Let $\psi$ be an MSO formula. If $\psi$ has infinitely many models of out-degree $1$, then there exist $\tilde{\Gamma}=\{\tilde{G}_{1},\tilde{G}_{2},\tilde{G}_{3}\}$ three $3$-graphs such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)\models\psi$ for all $\ell\in\mathbb{N}$, and $P_{1}(\tilde{G_{1}})\cap P_{2}(\tilde{G_{1}})\neq V(\tilde{G_{1}})$. Moreover, $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ has out-degree $1$ for all $\ell\in\mathbb{N}$.
@@
$\chi=\forall x,\exists y,(x\to y)\wedge(\forall z,z\neq y\implies\neg(x\to z)),$
such that $G\vDash\chi$ if and only if the graph $G$ has out-degree exactly $1$ (is a deterministic dynamics, or a functional graph).
Recall that $\Sigma_{k,\psi}$ denotes the set of equivalence classes of $\sim_{k,\psi}$ for $k$-graphs. It is finite for any given $k$ and $\psi$, hence $\Sigma_{k,\psi,\chi}=\Sigma_{k,\psi}\times\Sigma_{k,\chi}$ is finite as well. In the rest of this section $k=3$, so we drop it from the notation. When considering $\Sigma_{\psi,\chi}$-labeled trees, each node $v$ has an assocatied equivalence classe $\mathcal{C}(v)$, and we also define $\mathcal{C}^{\psi}(v)$ and $\mathcal{C}^{\chi}(v)$ for the equivalences classes for $\psi$ and $\chi$, respectively. Hence, for two nodes $v$ and $v^{\prime}$ of a $\Sigma_{\psi,\chi}$-labeled tree, we have $\mathcal{C}(v)=\mathcal{C}(v^{\prime})$ if and only if $\mathcal{C}^{\psi}(v)=\mathcal{C}^{\psi}(v^{\prime})$ and $\mathcal{C}^{\chi}(v)=\mathcal{C}^{\chi}(v^{\prime})$. We are now able to prove the proposition.
@@
This proof brings the ideas of [10] to the deterministic setting. Since $\psi$ has infinitely many deterministic models, we consider a family $(M_{i})_{i\in\mathbb{N}}$ of deterministic models and $(D_{i})_{i\in\mathbb{N}}$ a sequence of their $3$-tree-decompositions, such that each $D_{i}$ has at least one path of $\mathcal{N}$-length $i$ (which is possible by splitting nodes of degree more than three in the tree). Without loss of generality we assume that every $M_{i}$ has at least $3$ vertices, and that all the bags of $D_{i}$ are of size $3$.
We view the tree-decompositions $D_{i}$ as $\Sigma_{\psi,\chi}$-labeled trees. Let $s=|\Sigma_{\psi,\chi}|$. Any path of $\mathcal{N}$-length at least $s+1$ in a tree-decomposition contains two $\mathcal{N}$-different nodes with the same label. In particular, $D_{s+1}$ contains a path with two $\mathcal{N}$-different vertices $v$ and $v^{\prime}$, such that $\mathcal{C}(v)=\mathcal{C}(v^{\prime})$. Assume that $v$ is higher than $v^{\prime}$ in the tree, and recall that $\mathcal{S}(v)$ is the largest subtree of $D_{s+1}$ rooted in $v$. We define:
$T_{1}=(\mathcal{S}(v)\backslash\mathcal{S}(v^{\prime}))\cup\{v^{\prime}\},\qquad T_{2}=(D_{s+1}\backslash\mathcal{S}(v))\cup\{v\},\qquad T_{3}=\mathcal{S}(v^{\prime}),$
so that we have $T_{2}\oplus T_{1}\oplus T_{3}=D_{s+1}$ (choosing the bags of $v$ and $v^{\prime}$ as ports). Since $\mathcal{C}(v)=\mathcal{C}(v^{\prime})$, and in particular $\mathcal{C}^{\psi}(v)=\mathcal{C}^{\psi}(v^{\prime})$, by [10, Lemma 14] we deduce that $\Lambda^{\mathcal{T}}(2\cdot 1^{\ell}\cdot 3)$ is positive for $\psi$ for all $\ell\in\mathbb{N}$. Then, by Remark 2, there exist $\tilde{\Gamma}=\{\tilde{G}_{1},\tilde{G}_{2},\tilde{G}_{3}\}$ associated respectively to $T_{1},T_{2},T_{3}$ such that $\Delta^{\tilde{\Gamma}}(2\cdot 1^{\ell}\cdot 3)$ is a model of $\psi$ for all $\ell\in\mathbb{N}$ (remark that by definition of $\mathcal{N}$-difference, there exists at least one vertex in $G_{1}$ which is not both a primary port nor a secondary port, that is $P_{1}(\tilde{G_{1}})\cap P_{2}(\tilde{G_{1}})\neq V(G_{1})$, hence the sequence $(|\Delta^{\Gamma}(2\cdot 1^{\ell}\cdot 3)|)_{\ell\in\mathbb{N}}$ is strictly increasing). Analogously, we deduce that $\Delta^{\Gamma}(2\cdot 1^{\ell}\cdot 3)$ is a model of $\chi$ for all $\ell\in\mathbb{N}$, hence that all these graphs have out-degree $1$. ∎
Let $\Omega_{m}$ denote the deterministic saturating graph from Corollary 3, employed to reach Conditions 1–3 of Proposition 4, and the positivity if and only if the $\Delta^{\Gamma}$ gluing contains at least one copy of $G_{0}$.
@@
Let $\tilde{\Gamma}=\{\tilde{G}_{1},\tilde{G}_{2},\tilde{G}_{3}\}$ be given by Proposition 6. Consider $\tilde{G_{0}}=\Omega_{m}\sqcup\tilde{G_{1}}$, such that Conditions 2–3 of Proposition 4 (relative to the ports) are verified, and $\tilde{G_{0}}$ is still deterministic. Now, let $G_{1}=\bigoplus_{i=1}^{n}\tilde{G_{1}}$ with $n$ such that $|G_{1}|\geq|\tilde{G_{0}}|$, and let $G_{0}$ be $\tilde{G_{0}}$ with $|G_{1}|-|\tilde{G_{0}}|$ additionnal isolated fixed points. We have $|G_{0}|=|G_{1}|$ (Condition 1), and the proof is completed by taking $G_{2}=\tilde{G}_{2}$ and $G_{3}=\tilde{G}_{3}$. Indeed, thanks to the deterministic saturating graph $\Omega_{m}$ from Corollary 3, it holds that $\Delta^{\Gamma}(2\cdot w\cdot 3)$ is a model of $\psi$ if and only if $w\in\{0,1\}^{*}$ contains at least one letter 00.
It is now straightforward to build a polytime metareduction from $\Gamma=\{G_{0},G_{1},G_{2},G_{3}\}$, as in Proposition 4 from [10] (see Section 3.2). An out-degree $1$ graph on $N$ vertices is encoded as a circuit with $\log_{2}(N)$ input bits and $\log_{2}(N)$ output bits. ∎

### v2

@@ 3.3 A reduction from SAT to $\psi$-NAN-dynamics
The neutral is intending to “change nothing”, and is obtained by pumping (see Section 3.2). Again, we consider clique-decompositions instead of the graphs they generate.
Recall that a clique-decomposition corresponds to a unique graph, therefore we abuse some notations. We first introduce a convenient way to deal with the assignments of $S$, then state our main result, and present our construction in a separate statement.
Definition 1.
If $S$ is an instance of SAT with $s$ variables, then $\overline{S}$ is the word of length $2^{s}$ such that $\overline{S}_{i}$ is $1$ if $S(i)$ is false, and 0 if $S(i)$ is true (viewing the binary expansion of $i$ as a Boolean assignment for $S$). And $\smash{\overline{\overline{S}}}$ is the word $\overline{S}$ where we insert an additional $1$ in front of each symbol, hence of size $2^{s+1}$.
For example, let $S=(x_{1}\vee\neg x_{2})\wedge x_{3}$. Then $\overline{S}=11110010$, with $\overline{S}_{7}=0$ because the binary expansion of $7$ is $111$ and $(1\vee\neg 1)\wedge 1$ is true. And $\smash{\overline{\overline{S}}}=1111111110101110$.
When $\mathcal{C}$ is a marked clique-decomposition and $\mathcal{C}^{\prime}$ is the clique-decomposition of a saturating graph, then we denote $\mathcal{C}\oplus\mathcal{C^{\prime}}=\mathcal{C}\sqcup\mathcal{C^{\prime}}$ with the same marked node as $\mathcal{C}$.
Proposition 7.
Fix an MSO sentence $\psi$. Let $k$ be an integer and $\Gamma=\{\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ four $k$-clique-decompositions satisfying the following conditions:
  1. (i)
    $\mathcal{C}_{0}$ is the clique-decomposition of a saturating graph;
  2. (ii)
    $\mathcal{C}_{0}$ and $\mathcal{C}_{1}$ have the same number of constant nodes;
  3. (iii)
    $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are marked.
@@
Suppose that, for every word $w$ over alphabet $\{0,1\}$, we have $\Delta^{\Gamma}(2\cdot w\cdot 3)\models\psi$ if and only if $w$ contains letter 0. Then $\psi$-NAN-dynamics is $\mathsf{NP}$}-hard.
The reduction from a SAT instance $S$ must produce the circuit of a NAN whose transition digraph corresponds to a clique-decomposition obtained by pumping (Proposition 5), as illustrated on Figure 5. Apart from some rework of the clique-decomposition to ensure convenient properties on the recolor nodes (stated in Lemma 8 and obtained in the proof of Proposition 7), the construction presents no difficulty. It should also become clear why we interleave copies of $\mathcal{C}_{1}$. Given a $k$-clique-decomposition $\mathcal{C}$ with marked node $v$ and $C=\{1,\dots,k\}$, the composition of recolorings from $v$ to the root $r$ of $\mathcal{C}$ is simply the composition of functions $f:C\to C$ for all the recolorf nodes in the path from $v$ to $r$ (from the leaf to the root). From an initial color $c(v)$, it computes the color of node $v$ in the $k$-colored graph constructed by $\mathcal{C}$. A function $g:C\to C$ is idempotent when $g^{2}=g$.
[…]
Figure 5: Clique decomposition of $\Delta^{\Gamma}(2\cdot\smash{\overline{\overline{S}}}\cdot 3)$ in the statement of Lemma 8, for the word $\smash{\overline{\overline{S}}}=1111111110101110$ corresponding to the SAT instance $S=(x_{1}\vee\neg x_{2})\wedge x_{3}$. Observe that disjoint unions for $\mathcal{C}_{0}$ employ the join.
Lemma 8.
Let $S$ be an instance of SAT with $s$ variables, $k$ an integer and $\Gamma=\{\mathcal{C}_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ four $k$-clique-decompositions satisfying conditions (i)(iii) from Proposition 7. Let $C=\{1,\dots,k\}$ and $g_{1}:C\to C$ be the composition of recolorings from the marked node of $\mathcal{C}_{1}$ to its root. If $g_{1}$ is idempotent, then there exists a NAN $f$ of size $n$ represented by circuit $D$ that can be computed in polynomial time given $S$, such that $\Delta^{\Gamma}(2\cdot\smash{\overline{\overline{S}}}\cdot 3)$ is a clique-decomposition of $\mathcal{G}_{f}$ (the $\mathcal{C}_{i}$ are considered constant).
Proof.
Let $n_{1}$ be the number of constant nodes in $\mathcal{C}_{0}$ and $\mathcal{C}_{1}$ which are equal by Condition (ii), and $n_{2}$ (resp. $n_{3}$) be the number of constant nodes in $\mathcal{C}_{2}$ (resp. $\mathcal{C}_{3}$). We set $N=n_{2}+2^{s+1}\cdot n_{1}+n_{3}$, which is the number of constant nodes in $\Delta^{\Gamma}(2\cdot\smash{\overline{\overline{S}}}\cdot 3)$. Recall that constant nodes define the vertices of $k$-colored graphs, hence we identify them in our explanations. In terms of the circuit $D$, the vertex labels of $\mathcal{G}_{f}$ are $\{0,\dots,n_{2}+2^{s+1}\cdot n_{1}+n_{3}-1\}$: see Figure 6 for a picture of its organization. The initial segment corresponds to the (unique) copy of $\mathcal{C}_{2}$, followed by $2^{s+1}$ copies of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$, followed by the (unique) copy of $\mathcal{C}_{3}$.
[…]
Figure 6: Proof of Lemma 8, a representation of $\{0,\dots,n_{2}+2^{s+1}\cdot n_{1}+n_{3}-1\}$ as the set of vertex labels of $\mathcal{G}_{f}$ on NN configurations.
In order to encode the NAN $f$ whose dynamics $\mathcal{G}_{f}$ has clique-decomposition $\Delta^{\Gamma}(2\cdot\smash{\overline{\overline{S}}}\cdot 3)$, the circuit $D$ needs to decide whether there is an edge from $v$ to $v^{\prime}$. This depends on the set $M$ of their least common ancestor node joinM, and on their respective colors in the two children $k$-colored graphs. That is, the circuit needs to compute the result of recolor nodes upward. The idempotence condition makes this process finite and straightforward.
Given $v$ and $v^{\prime}$ two vertex labels in $\{0,\dots,N-1\}$, the circuit $D$ first computes the corresponding constant nodes in the clique-decomposition. It can belong to:
  • the copy of $\mathcal{C}_{2}$,
  • one of the $2^{s+1}$ copies of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$,
  • the copy of $\mathcal{C}_{3}$.
In order to compute the vertex label relative to a clique-decomposition from $\Gamma$ (on a constant number of bits), and to keep track of the index among the $2^{s+1}$ copies of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$ (on $s+2$ bits, to account for the unique copies of $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$), we define respectively:
[…]
$\gamma(v)$ $=\begin{cases}0&\text{ if }v<n_{2},\\ 1+\left\lfloor\frac{v-n_{2}}{n_{1}}\right\rfloor&\text{ if }n_{2}\leq v<n_{2}+2^{s+1}\cdot n_{1},\\ 2+2^{s+1}&\text{ if }n_{2}+2^{s+1}\cdot n_{1}\leq v.\end{cases}$ (2)
When $\gamma(v)\in\{2,4,6,\dots,2^{s+1}\}$, the circuit needs to evaluate $S$ on the Boolean assignment given by the $s$ first bits of $\gamma(v)$, i.e., $\frac{\gamma(v)}{2}$, to decide whether $v$ belongs to a copy of $\mathcal{C}_{0}$ (true) or $\mathcal{C}_{1}$ (false). In all other cases it is straightforward to identify the clique-decomposition from $\Gamma$ to which $v$ belongs.
In order to obtain the set $M$ of the node joinM that is the least common ancestor of $v$ and $v^{\prime}$, and to compute their respective colors in the two children $k$-colored graphs, we have the following cases.
  • If $\gamma(v)=\gamma(v^{\prime})$ then $v$ and $v^{\prime}$ belong to the same copy of a clique-decomposition from $\Gamma$, and their adjacency is decided in constant time from their relative constant nodes $\delta(v)$ and $\delta(v^{\prime})$ in that clique-decomposition.
  • Else, if $\gamma(v)\neq\gamma(v^{\prime})$ and at least one of them belongs to a copy of $\mathcal{C}_{0}$, then there is no arc from $v$ to $v^{\prime}$ because $M=\emptyset$ by the disjoint union of the saturating graph.
  • Else, if $|\gamma(v)-\gamma(v^{\prime})|\geq 3$, we present the case $\gamma(v)>\gamma(v^{\prime})$ (the other case is symmetric). Let $\mathcal{C}_{x}$ be the clique-decomposition of $v$ within $\Gamma$, and $\mathcal{C}_{y}$ be the clique-decomposition of $v^{\prime}$ within $\Gamma$ (we can assume $x,y\in\{1,2,3\}$). First, observe that the joinM node belongs to the clique-decomposition of $v$: it is the join of $v$ and the marked node of $\mathcal{C}_{x}$. Consequently the color of $v$ in its child $k$-colored graph depends only on $\mathcal{C}_{x}$ and is computed in constant time from $\delta(v)$. Second, observe that the path from $v^{\prime}$ to this joinM node contains at least one full traversal of a copy of $\mathcal{C}_{1}$ (here the interleaved copies of $\mathcal{C}_{1}$ are useful). Consequently, the color of $v^{\prime}$ in its child $k$-colored graph is computed from its initial color $c(v^{\prime})$ as a constant node in $\mathcal{C}_{y}$, by applying: the finitely many recolorings in the path to the root of $\mathcal{C}_{y}$; then once $g_{1}$, which accounts for any number of copies of $\mathcal{C}_{1}$ by idempotency; and lastly the finitely many recolorings from the marked node of $\mathcal{C}_{x}$ to the joinM node. From the eventual colors of $v,v^{\prime}$ and the set $M$, one decides whether the arc $(v,v^{\prime})$ exists.
  • Else, if $|\gamma(v)-\gamma(v^{\prime})|\leq 2$, then there are finitely many cases to check (each requiring finite explorations from $\delta(v)$ and $\delta(v^{\prime})$), depending on $\gamma(v),\gamma(v^{\prime})$, their respective clique-decompositions within $\Gamma$, and whether they are separated by a copy of $\mathcal{C}_{0}$ or by a copy of $\mathcal{C}_{1}$ (when $|\gamma(v)-\gamma(v^{\prime})|=2$).
The algorithm described above can immediately be implemented as a circuit $D$ produced in polynimial time, because the clique-decompositions $\mathcal{C}_{i}$ are considered constants, hence the only non-constant parts correspond to computing $\delta,\gamma$ and evaluating $S$ (at most twice). This concludes the proof. ∎
We are now equipped to prove Proposition 7, leading to Theorem 1.c. The former bridges the gap to obtain the idempotency condition, and the latter assembles our Propositions from the previous sections.
Proof of Proposition 7.
Given $\mathcal{C}_{1}$ from the statement, consider $g_{1}:C\to C$ the composition of recolorings from the marked node of $\mathcal{C}_{1}$ to its root, with $C=\{1,\dots,k\}$. Since each element of a finite semigroup has an idempotent power, there is $p$ such that $g_{1}^{p}:C\to C$ is idempotent. Defining $\mathcal{C}^{\prime}_{1}=\bigoplus_{i=1}^{p}\mathcal{C}_{1}$ and $\mathcal{C}^{\prime}_{0}=\bigoplus_{i=1}^{p}\mathcal{C}_{0}$, the set $\Gamma^{\prime}=\{\mathcal{C}^{\prime}_{0},\mathcal{C}^{\prime}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ fulfills all the conditions to apply Lemma 8. It is now an immediate consequence of Lemma 8: given an instance $S$ of SAT, compute in polynomial time the corresponding circuit $D$ of $f$ on $N$ configurations, then ask $\psi$-NAN-dynamics on it. ∎
Proof of Theorem 1.c.
Let $\psi$ denote a cw-non-trivial MSO sentence with $m$ quantifiers. By Proposition 3, the graph $\Omega_{m}$ is either forbidden or sufficient for $\psi$. Assume that $\Omega_{m}$ is sufficient for $\psi$ and prove that $\psi$-NAN-dynamics is $\mathsf{NP}$}-hard. (If it was forbidden, then consider $\neg\psi$ instead—this would yield that $\psi$-NAN-dynamics is $\mathsf{coNP}$}-hard.) Let $\mathcal{C}_{0}$ be a $k$-clique-decomposition of $\Omega_{m}$ (this may require a larger $k$ than the value given by cw-non-triviality, but $\Omega_{m}$ is constant in our context, so we can just consider their maximum). By Proposition 5 applied to $\neg\psi$, there exist a triple of $k$-clique-decompositions $\Gamma=\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ such that $\Delta^{\Gamma}(2\cdot 1^{n}\cdot 3)$ is a model of $\neg\psi$ for every integer $n$. Let $\mathcal{C}^{\prime}_{1}=\bigoplus_{i=1}^{n}\mathcal{C}_{1}$, where $n$ is the smallest integer such that the number of constant nodes in $\mathcal{C}^{\prime}_{1}$ is greater or equal to that of $\mathcal{C}_{0}$. Finally, let $\mathcal{C}^{\prime}_{0}$ denote $\mathcal{C}_{0}$ with enough isolated vertices added so that its number of constant nodes equals that of $\mathcal{C}^{\prime}_{1}$ (adding an isolated vertex to a clique-decomposition corresponds to adding a constant node under a join at the root). The quadruple $\Gamma^{\prime}=\{\mathcal{C}^{\prime}_{0},\mathcal{C}^{\prime}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ satisfies all the conditions of Proposition 7. This concludes the proof of the theorem. ∎
Remark that in the proof of Theorem 1.c above (in particular in the proof of Lemma 8 and its use in the proof of Proposition 7), we have no asumption on the number $N$ of vertices (configurations) in the transition graph we produce. If $N$ is not a power of $q$, then it is impossible to get a $q$-uniform NAN. We solve this issue in this work.
@@ 3.4 Contributions
  • A careful consideration in gluing (with $\oplus$) deterministic dynamics (graphs of out-degree 11 are handled through Hanf’s local lemma in [11]) in Section 4.
  • The addition of arithmetical considerations in Section 5, to restrict the constructions to $q$-uniform networks in Section 6.
@@ 4 Complexity lower bounds for deterministic networks
In this section, we extend Theorem 1.a for FO to MSO logics, or more accurately we transpose the proof technique of Theorem 1.c from non-deterministic NANs to deterministic ANs (whose transition digraphs have out-degree $1$). Recall that the non-triviality of formulas is defined relative to graphs of out-degree $1$ only. We prove the following part of Theorem 1.
@@
We follow the proof structure presented in Section 3, using the deterministic saturating graph from Corollary 4. The main difficulty consists in transposing Proposition 5 in order to obtain graphs of out-degree $1$ all along the way (i.e., to ensure that gluings preserve this property), and meet the required conditions for Proposition 7. We combine these into the following proposition (out-degree $1$ graphs have cliquewidth at most $6$  [^2] Any graph of out-degree $1$ has treewidth at most $2$ (its connected components are cycles with upward trees rooted on it), thus by [4] it has cliquewidth at most $6$.).
@@
Let $\psi$ be an MSO sentence and $k$ an integer. If $\psi$ has infinitely many models of out-degree $1$, then there exist clique-decompositions $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ such that $\mathcal{C}_{2}\oplus\mathcal{C}_{1}\oplus\dots\oplus\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ is a model of $\psi$ of out-degree $1$, whatever the number of occurrences of $\mathcal{C}_{1}$. Furthermore, $\mathcal{C}_{1}$ contains at least one constant node.
@@
$\chi=\left[\forall x,\exists y,(x\to y)\right]\wedge\left[\forall x,\forall y,\forall y^{\prime},\neg(x\to y)\vee\neg(x\to y^{\prime})\vee(y=y^{\prime})\right],$
such that $G\models\chi$ if and only if the graph $G$ has out-degree exactly $1$ (is a deterministic dynamics, or a functional graph).
Recall that $T_{m}$ denotes the set of realized types of quantifier rank $m$, which is finite for any given $k$ and $m$. In particular, if $m=\max(m^{\prime},2)$ where $m^{\prime}$ is the quantifier rank of $\psi$ and $2$ the quantifier rank of $\chi$ and $k=6$, then $T_{m}$ is finite. In the following, when considering a $k$-clique-decomposition $\mathcal{C}$, each node $v$ defines a maximal sub-clique-decomposition denoted $\mathcal{C}(v)$ (i.e., the clique-decomposition containing all the descendant of $v$) that is rooted in $v$, and the type of $\mathcal{C}(v)$ is written $t(v)$.
@@
This proof brings the ideas of [10] to the deterministic setting. Since $\psi$ has infinitely many deterministic models, and since there are finitely many types, there must be a type $t_{0}\in T_{m}$ which contains $\psi$ and $\chi$ and is realized by infinitely many graphs of cliquewidth at most $k$. Let $\mathcal{C}$ be a clique-decomposition of cliquewidth at most $k$ of height at least $|T_{m}|+1$, generating a graph of type $t_{0}$.
Any path of length at least $|T_{m}|+1$ in a clique-decomposition contains at least two different nodes $v$ and $v^{\prime}$ such that $t(v)=t(v^{\prime})$. Let $v$ and $v^{\prime}$ be two such nodes in a path of length at least $|T_{m}|+1$ in $\mathcal{C}$. We assume without loss of generality that $v^{\prime}$ is a descendant of $v$. Let $\mathcal{C}_{3}=\mathcal{C}(v^{\prime})$, $\mathcal{C}_{1}$ be $\mathcal{C}(v)$ where $\mathcal{C}(v^{\prime})$ has been removed and $v^{\prime}$ has been replaced by $\square$, and $\mathcal{C}_{2}$ be $\mathcal{C}$ where $\mathcal{C}(v)$ has been removed and $v$ has been replaced by $\square$. These definitions are such that $\mathcal{C}=\mathcal{C}_{2}\oplus\mathcal{C}_{1}\oplus\mathcal{C}_{3}$. Remark that $\mathcal{C}_{1}$ is not empty since $v\neq v^{\prime}$ (Remark 2).
Considering the map $\Lambda_{\mathcal{C}_{1}}$ from Lemma 6, since the type of the graph generated by $\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ is the same as the one of the graph generated by $\mathcal{C}_{3}$, we have $\Lambda_{\mathcal{C}_{1}}(t(v))=t(v)=t(v^{\prime})=\Lambda_{\mathcal{C}_{1}}(t(v^{\prime}))$. So the type of the graph generated by $\mathcal{C}_{1}\oplus...\oplus\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ is $t(v)$, whatever the number of repetitions of $\mathcal{C}_{1}$. Finally, since $\mathcal{C}=\mathcal{C}_{2}\oplus\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ has type $t_{0}$, we have $\Lambda_{\mathcal{C}_{2}}(t(v))=t_{0}$ and the type of $\mathcal{C}_{2}\oplus\mathcal{C}_{1}\oplus...\oplus\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ is always $t_{0}$, and the graph they are generating are all models of both $\psi$ and $\chi$. ∎
Let $\Omega_{m}$ denote the deterministic saturating graph from Corollary 4, employed to reach the conditions of Proposition 7.
@@
$\psi$ has an infinite number of deterministic models and deterministic counter-models of quantifier rank $q^{\prime}$. Let $q=\max(q^{\prime},2)$, and suppose that $\Omega_{q}\vDash\psi$. Since $\psi$ has infinitely many counter-models of out-degree $1$, let $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ be three clique-decompositions obtained by Proposition 9, such that $\mathcal{C}_{1}$ contains at least one constant node and $\mathcal{C}_{2}\oplus\mathcal{C}_{1}\oplus\dots\oplus\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ is a model of $\neg\psi$ of out-degree 1, whatever the number of occurrences of $\mathcal{C}_{1}$.
Let $\mathcal{C}^{\prime}_{1}=\bigoplus_{i=1}^{n}\mathcal{C}_{1}$, where $n$ is the smallest integer such that the number of constant nodes in $\mathcal{C}^{\prime}_{1}$ is greater or equal to that of $\mathcal{C}_{0}$. Finally, let $\mathcal{C}_{0}$ denote $\Omega_{m}$ with enough isolated vertices added so that its number of constant nodes equals that of $\mathcal{C}^{\prime}_{1}$ (adding an isolated vertex to a clique-decomposition corresponds to adding a constant node under a join at the root). The quadruple $\Gamma^{\prime}=\{\mathcal{C}_{0},\mathcal{C}^{\prime}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ satisfies all the conditions of Proposition 7. In the case where $\Omega_{m}\nvDash\psi$, we have $\Omega_{m}\vDash\neg\psi$, hence the same proof applies to $\neg\psi$. This concludes the proof of the theorem. ∎